# Riemannian Geometry of Optimal Rebalancing in Dynamic Weight Automated Market Makers

### 09 Mar 2026 v2

@@ A.2 Deferred Proof: Corollary 1 [Quadratic loss kernel]
$\log r=\sum_{i=1}^{N}w_{i}^{\mathrm{end}}\log\!\left(\frac{w_{i}^{\mathrm{start}}}{w_{i}^{\mathrm{end}}}\right).$ (A.9)
@@
$=-\sum_{i=1}^{N}(w_{i}+\Delta w_{i})\log\!\left(1+\frac{\Delta w_{i}}{w_{i}}\right).$ (A.10)
@@
$=-\underbrace{\sum_{i=1}^{N}\Delta w_{i}}_{=\,0}-\sum_{i=1}^{N}\frac{(\Delta w_{i})^{2}}{2w_{i}}+O(\Delta w^{3}).$ (A.11)
@@
$\log r\;\approx\;\sum_{i=1}^{N}\frac{(\Delta w_{i})^{2}}{2w_{i}},$ (A.12)
@@ Appendix C Deferred proof: Theorem 3 [sub-optimality bound]
and $R_{k}=\sum_{i}w_{i}^{(k-1)}\,h(u_{i}^{(k)})$ with $u_{i}^{(k)}=\Delta w_{i}^{(k)}/w_{i}^{(k-1)}$ and $h(u)=(1{+}u)\log(1{+}u)-u-u^{2}/2$. SLERP minimises $Q$, so $\nabla Q\big|_{\mathrm{SLERP}}=0$ and $\nabla C\big|_{\mathrm{SLERP}}=\nabla R\big|_{\mathrm{SLERP}}$.
@@
Each component of $\nabla R$ with respect to $w_{i}^{(k)}$ receives contributions from steps $k$ and k+1k{+}1. Since $h^{\prime}(u)=\log(1{+}u)-u$ satisfies $|h^{\prime}(u)|\leq u^{2}/(1-|u|)$ for $|u|<1$, and $|u_{i}^{(k)}|\leq 2\Omega/(f\epsilon)\leq 1/2$ for $f\geq 4\Omega/\epsilon$, each gradient component has magnitude
@@
The Hessian of $Q$ restricted to the simplex constraints is block-diagonal in the intermediate weight vectors. Each diagonal block receives contributions from two consecutive Fisher–Rao metrics $\mathrm{diag}(1/w_{i}^{(k-1)})$ and $\mathrm{diag}(1/w_{i}^{(k)})$ (the latter from step k+1k{+}1, where $w^{(k)}$ is the base point). On the simplex hyperplane $\{\sum_{i}v_{i}=0\}$, each Fisher–Rao metric has minimum eigenvalue $\geq 1/\max_{i}w_{i}\geq 1$, so each block satisfies $\lambda_{\min}\geq 2$. The per-step KL cost is proportional to $(1{+}u)\log(1{+}u)$, whose second derivative is 1/(1+u)1/(1{+}u), while the quadratic model has second derivative 11. The Hessian perturbation per block is therefore proportional to 1/(1+u)1=u/(1+u)1/(1{+}u)-1=-u/(1{+}u), which satisfies $|u/(1{+}u)|\leq|u|/(1-|u|)\leq 2|u|$ for $|u|\leq 1/2$. By the Step 1 bound $|u|\leq 2\Omega/(f\epsilon)$, the perturbation magnitude is at most $4\Omega/(f\epsilon)$ per block. For $f\geq 4\Omega/\epsilon$ (already required in Step 2) this is at most 11, so the Hessian of $C$ satisfies $\lambda_{\min}(\nabla^{2}C)\geq 2-1=1$ per block.
@@
$C_{\mathrm{SLERP}}-C_{*}\leq\frac{\|\nabla C\|^{2}}{2\,\lambda_{\min}(\nabla^{2}C)}\leq\frac{256\,N\,\Omega^{4}}{2\cdot f^{3}\,\epsilon^{4}}=\frac{128\,N}{f^{3}\,\epsilon^{4}}\,\Omega^{4},$
so $A=128\,N/\epsilon^{4}$ suffices. The relative sub-optimality is $A\,\Omega^{4}/f^{3}$ divided by the total SLERP cost $2\Omega^{2}/f$, giving $O(\Omega^{2}/f^{2})$. ∎
Appendix D Taylor agreement of interpolation methods
Write $w_{i}\equiv w_{i}^{\mathrm{start}}$ and $u_{i}=\Delta w_{i}/w_{i}$ where $\Delta w_{i}=w_{i}^{\mathrm{end}}-w_{i}^{\mathrm{start}}$. The simplex constraint gives $\sum_{i}w_{i}u_{i}=0$. Define $\sigma^{2}=\sum_{j}w_{j}u_{j}^{2}$. We expand all three methods to second order in $u$.
@@ Appendix D Taylor agreement of interpolation methods
$\cos\Omega=\sum_{j}w_{j}\sqrt{1+u_{j}}=1+\tfrac{1}{2}\underbrace{\sum_{j}w_{j}u_{j}}_{=\,0}-\tfrac{1}{8}\sigma^{2}+O(u^{3}),$
so $\Omega^{2}=2(1-\cos\Omega)+O(u^{4})=\sigma^{2}/4+O(u^{3})$. The SLERP coefficients satisfy, for small $\Omega$,
@@
where $\alpha=\sin((1\!-\!t)\Omega)/\sin\Omega$, $\beta=\sin(t\Omega)/\sin\Omega$. Using $(1\!-\!t)^{3}+t^{3}=1-3t(1\!-\!t)$, the sum evaluates to $\alpha+\beta=1+\frac{t(1-t)}{2}\Omega^{2}+O(u^{4})$. Squaring and substituting $\Omega^{2}=\sigma^{2}/4$:
$w_{i}^{\mathrm{SLERP}}(t)=w_{i}+t\,\Delta w_{i}+\frac{t(1-t)}{4}\,w_{i}\!\left(\sigma^{2}-u_{i}^{2}\right)+O(u^{3}).$ (D.13)
@@
The arithmetic component is $w_{i}^{\mathrm{AM}}(t)=w_{i}(1+tu_{i})$. The geometric component is $w_{i}^{\mathrm{GM}}(t)=w_{i}(1+u_{i})^{t}=w_{i}(1+tu_{i}+\frac{t(t-1)}{2}u_{i}^{2}+O(u^{3}))$. Their unnormalised sum is
@@
$Z=\sum_{j}(w_{j}^{\mathrm{AM}}+w_{j}^{\mathrm{GM}})=2+\tfrac{t(t-1)}{2}\sigma^{2}+O(u^{3}).$
@@
$\breve{w}_{i}(t)=w_{i}+t\,\Delta w_{i}+\frac{t(1-t)}{4}\,w_{i}\!\left(\sigma^{2}-u_{i}^{2}\right)+O(u^{3}),$ (D.14)
which matches Eq (D.13) exactly through $O(u^{2})$.
@@
The optimal 2-step midpoint of [1] is $\tilde{w}_{i}^{*}=w_{i}(1+u_{i})/W_{0}(e(1+u_{i}))$. Expanding $W_{0}(e(1+u))$ around $u=0$ (where $W_{0}(e)=1$) by implicit differentiation of $We^{W}=e(1+u)$:
@@
Then $\tilde{w}_{i}^{*}=w_{i}(1+u_{i})(1+u_{i}/2-3u_{i}^{2}/16)^{-1}=w_{i}(1+u_{i}/2-u_{i}^{2}/16+O(u^{3}))$. This requires renormalisation; the normalisation factor is $\sum_{j}\tilde{w}_{j}^{*}=1-\sigma^{2}/16+O(u^{3})$, giving
$\hat{w}_{i}^{\mathrm{LW}}=w_{i}+\tfrac{1}{2}\Delta w_{i}+\tfrac{1}{16}\,w_{i}(\sigma^{2}-u_{i}^{2})+O(u^{3}).$ (D.15)
Setting $t=1/2$ in Eq (D.13) gives the same second-order term $\frac{(1/2)(1/2)}{4}=\frac{1}{16}$, confirming agreement through $O(u^{2})$.
@@
Appendix E Why (AM+GM)/normalise works: the $\alpha$-family of geodesics
The midpoint equivalence (Theorem 2) follows from the binomial identity, but it has a geometric interpretation in terms of a family of “straight lines” on the probability simplex [7].
@@ Appendix E Why (AM+GM)/normalise works: the $\alpha$-family of geodesics
  • $\alpha=-1$ (arithmetic/linear interpolation). Geodesics are straight lines in the ordinary simplex coordinates $w_{i}$. The midpoint is $w_{i}^{\mathrm{mid}}=(w_{i}^{\mathrm{start}}+w_{i}^{\mathrm{end}})/2$. This is the interpolation used by liquidity bootstrapping pools and by TFMM when weights change linearly [2].
@@
$\left(\sqrt{a}+\sqrt{b}\right)^{2}=\underbrace{(a+b)}_{\text{AM term}}+\underbrace{2\sqrt{ab}}_{\text{GM term}}$ (E.16)
@@
For general $t\neq 1/2$, the three $\alpha$-geodesics diverge, and the (AM+GM)/normalise trajectory, which averages the $\alpha=\pm 1$ trajectories pointwise, no longer lies exactly on the $\alpha=0$ geodesic. The third-order agreement reported in Table 1 reflects the smooth dependence of geodesics on $\alpha$: near any base point, geodesics of the different connections share the same tangent vector and second-order curvature, differing only at third order.
E.1 Differential-geometric context
@@
The $\alpha=0$ connection is the Levi-Civita connection of the Fisher–Rao metric: the unique torsion-free, metric-compatible connection. Its geodesics solve the standard Euler–Lagrange equation for arc-length minimisation, which is how the SLERP optimality result (Corollary 2) is derived, without ever computing Christoffel symbols. The Hellinger embedding sidesteps the Christoffel computation entirely: it maps the simplex isometrically onto a sphere, where the geodesics are known to be great circles. (This is analogous to solving geodesic problems by calculus of variations in a convenient coordinate system, rather than by the geodesic equation $\ddot{x}^{\mu}+\Gamma^{\mu}_{\nu\rho}\dot{x}^{\nu}\dot{x}^{\rho}=0$ directly.)
@@
This dual flatness gives rise to a canonical divergence: the unique divergence function compatible with the dually-flat structure. On the probability simplex, the canonical divergence is the KL divergence [7]. The fact that arbitrage cost equals the KL divergence (Theorem 1) means rebalancing cost is the canonical divergence of the simplex.
@@
Appendix F Explicit computation: SLERP vs Lambert W for $N=2$

### 31 Mar 2026 v3

@@ A.2. Deferred Proof: Corollary 3 [Quadratic loss kernel]
(A.9) $\log r=\sum_{i=1}^{N}w_{i}^{\mathrm{end}}\log\!\left(\frac{w_{i}^{\mathrm{start}}}{w_{i}^{\mathrm{end}}}\right).$
@@
(A.10) $=-\sum_{i=1}^{N}(w_{i}+\Delta w_{i})\log\!\left(1+\frac{\Delta w_{i}}{w_{i}}\right).$
@@
(A.11) $=-\underbrace{\sum_{i=1}^{N}\Delta w_{i}}_{=\,0}-\sum_{i=1}^{N}\frac{(\Delta w_{i})^{2}}{2w_{i}}+O(\Delta w^{3}).$
@@
(A.12) $\log r\;\approx\;\sum_{i=1}^{N}\frac{(\Delta w_{i})^{2}}{2w_{i}},$
@@ Appendix C Deferred proof: Theorem 9 [sub-optimality bound]
and $R_{k}=\sum_{i}w_{i}^{(k-1)}\,\tilde{h}(u_{i}^{(k)})$ with $u_{i}^{(k)}=\Delta w_{i}^{(k)}/w_{i}^{(k-1)}$ and $\tilde{h}(u)=(1{+}u)\log(1{+}u)-u-u^{2}/2$ (the KL remainder beyond the quadratic; note this differs from the full per-component KL function $h(u)=(1{+}u)\log(1{+}u)-u$ used in Appendix M). SLERP minimises $Q$, so $\nabla Q\big|_{\mathrm{SLERP}}=0$ and $\nabla C\big|_{\mathrm{SLERP}}=\nabla R\big|_{\mathrm{SLERP}}$.
@@
Each component of $\nabla R$ with respect to $w_{i}^{(k)}$ receives contributions from steps $k$ and k+1k{+}1. Since $\tilde{h}^{\prime}(u)=\log(1{+}u)-u$ satisfies $|\tilde{h}^{\prime}(u)|\leq u^{2}/(1-|u|)$ for $|u|<1$, and $|u_{i}^{(k)}|\leq 2\Omega/(f\epsilon)\leq 1/2$ for $f\geq 4\Omega/\epsilon$, each gradient component has magnitude
@@
The Hessian of $C$ restricted to the simplex constraints is block-diagonal in the intermediate weight vectors. Each block receives contributions from two consecutive KL terms. For the first, $D_{\mathrm{KL}}(\boldsymbol{\mathbf{w}}^{(k)}\|\boldsymbol{\mathbf{w}}^{(k-1)})$, differentiating the first-argument contribution $w_{i}^{(k)}\log(w_{i}^{(k)}/w_{i}^{(k-1)})$ twice gives $\partial^{2}/\partial(w_{i}^{(k)})^{2}=1/w_{i}^{(k)}\geq 1$ (since $w_{i}\leq 1$). For the second, $D_{\mathrm{KL}}(\boldsymbol{\mathbf{w}}^{(k+1)}\|\boldsymbol{\mathbf{w}}^{(k)})$, the second-argument contribution $w_{i}^{(k+1)}\log w_{i}^{(k)}$ gives $\partial^{2}/\partial(w_{i}^{(k)})^{2}=w_{i}^{(k+1)}/(w_{i}^{(k)})^{2}\geq 1/2$ (using $w_{i}^{(k+1)}/w_{i}^{(k)}=1+u_{i}\geq 1/2$ for $|u|\leq 1/2$, and $w_{i}^{(k)}\leq 1$). On the simplex hyperplane, each block therefore has $\lambda_{\min}\geq 1+1/2=3/2$.
@@
$C_{\mathrm{SLERP}}-C_{*}\leq\frac{\|\nabla C\|^{2}}{2\,\lambda_{\min}(\nabla^{2}C)}\leq\frac{1}{3}\cdot\frac{256\,N\,\Omega^{4}}{f^{3}\,\epsilon^{4}}=\frac{256\,N}{3\,f^{3}\,\epsilon^{4}}\,\Omega^{4},$
so $A=256N/(3\epsilon^{4})\leq 86\,N/\epsilon^{4}$ suffices (using $\lambda_{\min}\geq 3/2$). The relative sub-optimality is $A\,\Omega^{4}/f^{3}$ divided by the total SLERP cost $2\Omega^{2}/f$, giving $O(\Omega^{2}/f^{2})$. ∎
Appendix D Deferred proof: modified geodesic and two-token pendulum
See 1
Proof.
The cost functional has Lagrangian $L=\tfrac{1}{2f}g_{ij}\dot{w}^{i}\dot{w}^{j}+T\ell(\boldsymbol{\mathbf{w}})$, where $g_{ij}=\delta_{ij}/w_{i}$ is the Fisher–Rao metric. The Euler–Lagrange equation gives:
$\frac{1}{f}\frac{d}{ds}\!\left(\frac{\dot{w}^{i}}{w_{i}}\right)=T\frac{\partial\ell}{\partial w^{i}}+\frac{1}{2f}\frac{(\dot{w}^{i})^{2}}{w_{i}^{2}}$
Expanding the left-hand side:
$\frac{1}{f}\!\left(\frac{\ddot{w}^{i}}{w_{i}}-\frac{(\dot{w}^{i})^{2}}{w_{i}^{2}}\right)=T\frac{\partial\ell}{\partial w^{i}}+\frac{1}{2f}\frac{(\dot{w}^{i})^{2}}{w_{i}^{2}}$
Multiplying through by $w_{i}$ and recognising that $\Gamma^{i}_{jk}=-\delta_{ij}\delta_{ik}/(2w_{i})$ for the Fisher–Rao metric on the simplex, so that $\Gamma^{i}_{jk}\dot{w}^{j}\dot{w}^{k}=-(\dot{w}^{i})^{2}/(2w_{i})$, the kinetic terms combine to give
$\frac{1}{f}\!\left(\ddot{w}^{i}+\Gamma^{i}_{jk}\dot{w}^{j}\dot{w}^{k}\right)=T\,w_{i}\frac{\partial\ell}{\partial w^{i}}=T\,g^{ij}\frac{\partial\ell}{\partial w^{j}}$
where $g^{ij}=w_{i}\,\delta_{ij}$ is the inverse Fisher–Rao metric. ∎
See 2
Proof.
For $N=2$ with weights (w,1w)(w,1-w), $\sigma_{1}=\sigma$, $\sigma_{2}=0$, the Hellinger coordinate $\theta=\arcsin(\sqrt{w})$ gives $w=\sin^{2}\theta$ and $1-w=\cos^{2}\theta$. The Fisher–Rao kinetic term becomes $\frac{1}{2f}[(\dot{w})^{2}/w+(\dot{w})^{2}/(1-w)]=\frac{2}{f}\dot{\theta}^{2}$ (using $\dot{w}=2\sin\theta\cos\theta\,\dot{\theta}=\sin(2\theta)\dot{\theta}$). The LVR potential is $\ell=\frac{\sigma^{2}}{2}w(1-w)=\frac{\sigma^{2}}{8}\sin^{2}(2\theta)$. The Lagrangian is $L=\frac{2}{f}\dot{\theta}^{2}+T\frac{\sigma^{2}}{8}\sin^{2}(2\theta)$, and the Euler–Lagrange equation gives $\frac{4}{f}\ddot{\theta}=T\frac{\sigma^{2}}{4}\sin(4\theta)$, which simplifies to Eq (24). ∎
Appendix E Taylor agreement of interpolation methods
Write $w_{i}\equiv w_{i}^{\mathrm{start}}$ and $u_{i}=\Delta w_{i}/w_{i}$ where $\Delta w_{i}=w_{i}^{\mathrm{end}}-w_{i}^{\mathrm{start}}$. The simplex constraint gives $\sum_{i}w_{i}u_{i}=0$. Define $s^{2}=\sum_{j}w_{j}u_{j}^{2}$. We expand all three methods to second order in $u$.
@@ Appendix E Taylor agreement of interpolation methods
$\cos\Omega=\sum_{j}w_{j}\sqrt{1+u_{j}}=1+\tfrac{1}{2}\underbrace{\sum_{j}w_{j}u_{j}}_{=\,0}-\tfrac{1}{8}s^{2}+O(u^{3}),$
so $\Omega^{2}=2(1-\cos\Omega)+O(u^{4})=s^{2}/4+O(u^{3})$. The SLERP coefficients satisfy, for small $\Omega$,
@@
where $\alpha=\sin((1\!-\!t)\Omega)/\sin\Omega$, $\beta=\sin(t\Omega)/\sin\Omega$. Using $(1\!-\!t)^{3}+t^{3}=1-3t(1\!-\!t)$, the sum evaluates to $\alpha+\beta=1+\frac{t(1-t)}{2}\Omega^{2}+O(u^{4})$. [^10] This comes from expanding $\sin((1-t)\Omega)/\sin\Omega+\sin(t\Omega)/\sin\Omega$ to order $\Omega^{2}$. Squaring and substituting $\Omega^{2}=s^{2}/4$:
(E.13) $w_{i}^{\mathrm{SLERP}}(t)=w_{i}+t\,\Delta w_{i}+\frac{t(1-t)}{4}\,w_{i}\!\left(s^{2}-u_{i}^{2}\right)+O(u^{3}).$
@@
The arithmetic component (Eq (3)) is $w_{i}^{\mathrm{AM}}(t)=w_{i}(1+tu_{i})$. The geometric component (Eq (4)) is $w_{i}^{\mathrm{GM}}(t)=w_{i}(1+u_{i})^{t}=w_{i}(1+tu_{i}+\frac{t(t-1)}{2}u_{i}^{2}+O(u^{3}))$. Their unnormalised sum is
@@
$Z=\sum_{j}(w_{j}^{\mathrm{AM}}+w_{j}^{\mathrm{GM}})=2+\tfrac{t(t-1)}{2}s^{2}+O(u^{3}).$
@@
(E.14) $\breve{w}_{i}(t)=w_{i}+t\,\Delta w_{i}+\frac{t(1-t)}{4}\,w_{i}\!\left(s^{2}-u_{i}^{2}\right)+O(u^{3}),$
which matches Eq (E.13) exactly through $O(u^{2})$.
@@
The optimal 2-step midpoint of [Willetts and Harrington, 2024a] is $\tilde{w}_{i}^{*}=w_{i}(1+u_{i})/W_{0}(e(1+u_{i}))$. Expanding $W_{0}(e(1+u))$ around $u=0$ (where $W_{0}(e)=1$) by implicit differentiation of $We^{W}=e(1+u)$:
@@
Then $\tilde{w}_{i}^{*}=w_{i}(1+u_{i})(1+u_{i}/2-3u_{i}^{2}/16)^{-1}=w_{i}(1+u_{i}/2-u_{i}^{2}/16+O(u^{3}))$. This requires renormalisation; the normalisation factor is $\sum_{j}\tilde{w}_{j}^{*}=1-s^{2}/16+O(u^{3})$, giving
(E.15) $\hat{w}_{i}^{\mathrm{LW}}=w_{i}+\tfrac{1}{2}\Delta w_{i}+\tfrac{1}{16}\,w_{i}(s^{2}-u_{i}^{2})+O(u^{3}).$
Setting $t=1/2$ in Eq (E.13) gives the same second-order term $\tfrac{(1/2)(1/2)}{4}=\tfrac{1}{16}$, confirming agreement through $O(u^{2})$.
@@
Appendix F Why (AM+GM)/normalise works: the $\alpha$-family of geodesics
The midpoint equivalence (Theorem 6) follows from the binomial identity, but it has a geometric interpretation in terms of a family of “straight lines” on the probability simplex [Amari, 1985].
@@ Appendix F Why (AM+GM)/normalise works: the $\alpha$-family of geodesics
  • $\alpha=-1$ (arithmetic/linear interpolation). Geodesics are straight lines in the ordinary simplex coordinates $w_{i}$. The midpoint is $w_{i}^{\mathrm{mid}}=(w_{i}^{\mathrm{start}}+w_{i}^{\mathrm{end}})/2$. This is the interpolation used by liquidity bootstrapping pools and by TFMM when weights change linearly [team, 2023a].
@@
(F.16) $\left(\sqrt{a}+\sqrt{b}\right)^{2}=\underbrace{(a+b)}_{\text{AM term}}+\underbrace{2\sqrt{ab}}_{\text{GM term}}$
@@
For general $t\neq 1/2$, the three $\alpha$-geodesics diverge. The (AM+GM)/normalise trajectory averages the $\alpha=\pm 1$ trajectories pointwise and no longer lies exactly on the $\alpha=0$ geodesic. The third-order agreement reported in Table 1 reflects the smooth dependence of geodesics on $\alpha$: near any base point, geodesics of the different connections share the same tangent vector and second-order curvature, differing only at third order.
F.1. Differential-geometric context
@@
The $\alpha=0$ connection is the Levi-Civita connection of the Fisher–Rao metric [Hobson et al., 2006]: the unique torsion-free, metric-compatible connection. Its geodesics solve the standard Euler–Lagrange equation for arc-length minimisation, which is how the SLERP optimality result (Corollary 5) is derived, without ever computing Christoffel symbols. The Hellinger embedding sidesteps the Christoffel computation entirely: it maps the simplex isometrically onto a sphere, where the geodesics are known to be great circles. (This is analogous to solving geodesic problems by calculus of variations in a convenient coordinate system, rather than by the geodesic equation $\ddot{x}^{\mu}+\Gamma^{\mu}_{\nu\rho}\dot{x}^{\nu}\dot{x}^{\rho}=0$ directly.)
@@
This dual flatness gives rise to a canonical divergence: the unique divergence function compatible with the dually-flat structure [Eguchi, 1983]. On the probability simplex, the canonical divergence is the KL divergence [Amari, 1985]. The fact that arbitrage cost equals the KL divergence (Theorem 2) means rebalancing cost is the canonical divergence of the simplex.
@@
Appendix G Explicit computation: SLERP vs Lambert W for $N=2$